# Theoretical Bounds for Stable In-Context Learning

### v1

@@ C.1 Martingale-difference setting
is a matrix martingale-difference sequence. By applying the matrix Azuma–Hoeffding inequality (Tropp 2011; Hsu et al. 2012), we obtain concentration of $\hat{\Sigma}_{K}=\tfrac{1}{K}\sum_{t=1}^{K}\phi_{t}(x)\phi_{t}(x)^{\top}$ around its expectation, with variance proxy and tail probability essentially the same as in the i.i.d. case, up to a constant factor.
Lower bound under martingale differences
@@
C.2 Mixing sequences
Alternatively, suppose $\{\phi_{t}(x)\}$ is $\alpha$-mixing with coefficient $\alpha(s)\leq\rho^{s}$ for some $\rho\in(0,1)$. Following Paulin (2015) and Merlevède et al. (2009), matrix Bernstein inequalities for mixing sequences yield concentration with an effective sample size
@@ C.2 Mixing sequences
Lower bound under mixing
@@
$K\;\gtrsim\;(1+c_{\rho})\cdot\frac{\sigma^{2}\,\|\Sigma\|}{(\lambda_{\min}(\Sigma)-\delta)^{2}}\,\log\!\frac{d}{\xi}.$
C.3 Discussion
@@
Appendix D Extensions Beyond the Nominal Assumption
In the main paper we derived our non-asymptotic lower bound under the nominal assumption (A1), namely that representations $\phi(x)$ are fixed, i.i.d., and sub-Gaussian with covariance $\Sigma$. Here we provide extensions under weaker assumptions (A2)–(A4), complementing Appendix B and Appendix C. Overall, these results show that our theory is robust: the sufficient prompt length bound degrades smoothly under moderate violations of (A1), rather than failing abruptly.
D.1 Representation Drift (A2)
As discussed in Appendix B, if the representation depends on the context index $t$, modeled as $\phi_{t}(x)=\phi(x)+\Delta_{t}(x)$ with $\|\Delta_{t}(x)\|_{2}\leq\varepsilon$, then eigenvalue perturbation bounds (Weyl’s inequality) imply that $\lambda_{\min}(\Sigma_{t})$ is reduced by at most $O(\varepsilon)$. Combining this with matrix Bernstein yields a lower bound of the form
@@ D.1 Representation Drift (A2)
D.2 Relaxed Tail Assumptions (A3)
If $\phi(x)$ has heavier tails than sub-Gaussian—e.g., only sub-exponential or bounded fourth moments—then truncation and Bernstein’s inequality (Vershynin, 2018) give concentration at a slightly slower rate. In this case, the sufficient prompt length becomes
@@ D.2 Relaxed Tail Assumptions (A3)
D.3 Weak Dependence (A4)
When features are not independent across $t$, we can use concentration for martingale-difference sequences (Tropp, 2011) or $\alpha$-mixing processes (Paulin, 2015; Merlevède et al., 2009). Both cases yield an effective sample size of the form $K_{\mathrm{eff}}=K/(1+c_{\rho})$, where $c_{\rho}$ is a dependence factor. Thus, weak dependence inflates the sufficient prompt length by a multiplicative constant, but does not alter the overall scaling.
D.4 Summary of Extensions
@@
(A1) Nominal Fixed sub-Gaussian features with covariance $\Sigma$ $K\;\geq\;K_{\mathrm{nominal}}=O\!\left(\tfrac{\sigma^{2}\|\Sigma\|}{(\lambda_{\min}(\Sigma)-\delta)^{2}}\log\!\tfrac{d}{\xi}\right)$
@@
Table 4: Diagnostics of representation assumptions on GPT-2 small. Reported values are averages across 500 prompts with 95% bootstrap confidence intervals in parentheses.
@@
The empirical kurtosis values are close to the Gaussian baseline (3.0) for short prompts but increase with $K$, suggesting heavier tails at longer contexts, consistent with Assumption (A3). The drift statistic $D(K)$ grows gradually, supporting Assumption (A2) with small but non-negligible $\varepsilon$. The spectral statistic $\lambda_{\min}(\hat{\Sigma}_{K})$ decreases as $K$ increases, and its trend correlates with the empirical knee-point in ICL accuracy. Bootstrap confidence intervals confirm that these trends are statistically significant. Overall, the diagnostics support our claim that the nominal assumption (A1) is a reasonable approximation in practice, while the relaxed extensions (A2)–(A4) capture the observed deviations.
Appendix E Synthetic Stress Tests
To evaluate the robustness of our theoretical bound under controlled violations of Assumption (A1), we conducted synthetic stress tests along three axes: (i) bounded drift in representations (A2), (ii) heavy-tailed features (A3), and (iii) temporal dependence (A4). For each setting we generated synthetic features and tasks as described in Section G.8, measured empirical knee-points $K_{\text{emp}}$ of in-context learning performance, and compared them against the theoretical lower bound $K_{\mathrm{nominal}}$.
E.1 Bounded Drift
@@
E.2 Heavy-tailed Features
@@
Heavier tails inflate the required prompt length by a multiplicative constant, but truncation or normalization substantially reduces the inflation. This behavior aligns with the extended bound in Appendix D.
E.3 Temporal Dependence
@@
Empirical knee-points scale approximately as $(1+c_{\rho})\cdot K_{\mathrm{nominal}}$, consistent with our dependence extension in Appendix C.
E.4 Summary
@@
Appendix F Interpretability of the Proxy
Predictive variance control via $\lambda_{\min}$
@@ Appendix F Interpretability of the Proxy
See F.1 for details; it follows by the matrix inequality $A^{-1}(K\widehat{\Sigma}_{K})A^{-1}\preceq A^{-1}$ with $A=K\widehat{\Sigma}_{K}+\lambda I$ and the Rayleigh quotient bound $\mathbf{v}^{\top}A^{-1}\mathbf{v}\leq\|\mathbf{v}\|_{2}^{2}/\lambda_{\min}(A)$. ∎
@@
Proposition F (Appendix F) shows that a lower bound on $\lambda_{\min}(\hat{\Sigma}_{K})$ suffices to control predictive variance of in-context regression, providing a mechanistic rationale for why it correlates with stability.
@@
F.1 Proof sketch of Proposition
@@ Conclusion.
which establishes Proposition F. $\square$
@@
Appendix G Ablation Studies
@@
[…]
G.1 ratio Sensitivity
@@
G.2 Effective Rank Definition
@@ G.3 Sub-Gaussian Parameter
G.4 Covariance Norm
@@ G.5 Label Tokenization
G.6 Prompt Truncation Policy
@@
G.7 Theory Constant Scaling
@@
G.8 Synthetic Data Experiments
@@ G.9 Retrieval Strategy
Interpretation. Retrieval reshapes the empirical covariance around the test point, preserving a larger effective $\lambda_{\min}$ and tightening $K^{*}$.
@@ G.10 Model Scale
Appendix H Theoretical Constants: Sources, Mappings, and Practical Calibration
H.1 List of Constants and Quick Reference
@@
Symbol Type Occurrence Recommended Range Reference
2, 2/32,\,2/3 Structural Matrix Bernstein tail bound Fixed, non-tunable Tropp (2012, Thm. 1.4)
$c_{1}$ Absolute Fourth-moment bound $[1.5,6]$, recommend 3344 Vershynin (2018, §3.4)
$c_{2}$ Absolute Covariance deviation bound $[2,4]$, recommend 33 Koltchinskii–Lounici (2017, Thm. 3.1)
$c^{\prime},\,c_{3}$ Engineering Sample-size formulae $\geq 8$, typically 10101515 This work (two-stage scheme)
H.2 Theoretical Origins and Geometric Interpretation
@@
For the sum of independent centered random matrices $S=\sum Y_{k}$, Tropp (2012) shows
@@
Covariance operator deviation constant $c_{2}$.
Koltchinskii & Lounici (2017, Thm. 3.1) proved
@@
H.3 Coupling with Problem Parameters
@@
H.4 Three-Step Calibration Procedure

### v2

@@ E.1 Martingale-difference setting
is a matrix martingale-difference sequence. By applying the matrix Azuma–Hoeffding inequality, we obtain concentration of $\hat{\Sigma}_{K}=\tfrac{1}{K}\sum_{t=1}^{K}\phi_{t}(x)\phi_{t}(x)^{\top}$ around its expectation, with variance proxy and tail probability essentially the same as in the i.i.d. case, up to a constant factor.
Proposition E.1 (Lower bound under martingale differences).
@@
E.2 Mixing sequences
Alternatively, suppose $\{\phi_{t}(x)\}$ is $\alpha$-mixing with coefficient $\alpha(s)\leq\rho^{s}$ for some $\rho\in(0,1)$. Matrix Bernstein inequalities for mixing sequences yield concentration with an effective sample size
@@ E.2 Mixing sequences
Corollary E.2 (Lower bound under mixing).
@@
$K\;\gtrsim\;(1+c_{\rho})\cdot\frac{\sigma^{2}\,\|\Sigma\|}{(\lambda_{\min}(\Sigma)-\delta)^{2}}\,\log\!\frac{d}{\xi},$
where $c_{\rho}=O(1/(1-\rho))$ quantifies the strength of temporal dependence.
Interpretation.
Kurtosis stays near the Gaussian baseline (3.0) for shorter prompts but increases with $K$, indicating heavier tails at longer contexts and motivating the heavy-tail extension (A3). The drift statistic $D(K)$ (defined as the average $\ell_{2}$ deviation between contextualized and context-free representations) grows gradually with $K$, suggesting that representation drift accumulates with longer prompts, consistent with (A2).
A potentially counter-intuitive observation is that $\lambda_{\min}(\widehat{\Sigma}_{K})$ can decrease as $K$ increases. This does not contradict the nominal i.i.d. picture for two reasons. First, $\lambda_{\min}$ of a finite-sample second-moment estimator is not monotone in $K$ in general: adding new rank-one terms can increase anisotropy and therefore reduce the smallest eigenvalue. A simple example is when early samples span multiple directions but additional samples concentrate on a dominant direction, shifting mass toward the top eigenvectors.
Second (and more importantly in our setting), demonstrations are typically query-conditioned due to retrieval. As $K$ increases, newly appended demonstrations are often increasingly redundant with earlier ones (e.g., lower diversity among nearest neighbors), which can reduce the effective rank and the empirical spectral coverage. Therefore a downward trend in $\lambda_{\min}(\widehat{\Sigma}_{K})$ is an empirical signature of non-i.i.d. selection and mild dependence/drift, which is precisely why the robustness extensions (A2)–(A4) are relevant. Notably, the observed trend correlates with the empirical knee-point in ICL accuracy, supporting the use of spectral coverage as a stability diagnostic.
E.3 Discussion
@@
Appendix F Extensions Beyond the Nominal Assumption
In the main paper we derived our non-asymptotic lower bound under the nominal assumption (A1), namely that representations $\phi(x)$ are fixed, i.i.d., and sub-Gaussian with covariance $\Sigma$. Here we provide extensions under weaker assumptions (A2)–(A4), complementing Appendix D and Appendix E. Overall, these results show that our theory is robust: the sufficient prompt length bound degrades smoothly under moderate violations of (A1), rather than failing abruptly.
F.1 Representation Drift (A2)
As discussed in Appendix D, if the representation depends on the context index $t$, modeled as $\phi_{t}(x)=\phi(x)+\Delta_{t}(x)$ with $\|\Delta_{t}(x)\|_{2}\leq\varepsilon$, then eigenvalue perturbation bounds (Weyl’s inequality) imply that $\lambda_{\min}(\Sigma_{t})$ is reduced by at most $O(\varepsilon)$. Combining this with matrix Bernstein yields a lower bound of the form
@@ F.1 Representation Drift (A2)
F.2 Relaxed Tail Assumptions (A3)
If $\phi(x)$ has heavier tails than sub-Gaussian e.g., only sub-exponential or bounded fourth moments then truncation and Bernstein’s inequality give concentration at a slightly slower rate. In this case, the sufficient prompt length becomes
@@ F.2 Relaxed Tail Assumptions (A3)
F.3 Weak Dependence (A4)
When features are not independent across $t$, we can use concentration for martingale-difference sequences or $\alpha$-mixing processes. Both cases yield an effective sample size of the form $K_{\mathrm{eff}}=K/(1+c_{\rho})$, where $c_{\rho}$ is a dependence factor. Thus, weak dependence inflates the sufficient prompt length by a multiplicative constant, but does not alter the overall scaling.
F.4 Summary of Extensions
@@
(A1) Nominal Fixed sub-Gaussian features with second-moment matrix $\Sigma$ $K\;\geq\;K_{\mathrm{nominal}}=O\!\left(\tfrac{\sigma^{2}\|\Sigma\|}{(\lambda_{\min}(\Sigma)-\delta)^{2}}\log\!\tfrac{d}{\xi}\right)$
@@
Table 5: Diagnostics of nominal assumptions using encoder embeddings (e.g., all-mpnet-base-v2). Reported values are averages across 500 prompts with 95% bootstrap confidence intervals in parentheses.
@@
Kurtosis stays near the Gaussian baseline (3.0) for shorter prompts but increases with $K$, indicating heavier tails at longer contexts, consistent with Assumption (A3). The drift statistic $D(K)$, defined as the average $\ell_{2}$-norm difference between contextualized and context-free representations, grows gradually with $K$ (from 0.05 at $K=16$ to 0.21 at $K=128$). This progressive increase suggests that representation drift accumulates as more demonstrations are included, supporting Assumption (A2) with small but non-negligible drift. Crucially, the smooth growth of $D(K)$ implies that the additional sample size required to compensate for drift increases polynomially with the drift magnitude $\varepsilon$ (as shown in Theorem D.3), providing practitioners with a predictable relationship between observed drift and prompt length requirements. The spectral statistic $\lambda_{\min}(\hat{\Sigma}_{K})$ decreases as $K$ increases, and its trend correlates with the empirical knee-point in ICL accuracy. Overall, these diagnostics support the claim that nominal assumptions are useful approximations in practice, while the relaxed extensions (A2)–(A4) capture systematic deviations.
Appendix G Synthetic Stress Tests
To evaluate the robustness of our theoretical bound under controlled violations of Assumption (A1), we conducted synthetic stress tests along three axes: (i) bounded drift in representations (A2), (ii) heavy-tailed features (A3), and (iii) temporal dependence (A4). For each setting we generated synthetic features and tasks as described in Appendix G (Section G.1), measured empirical knee-points $K_{\text{emp}}$ of in-context learning performance, and compared them against the theoretical lower bound $K_{\mathrm{nominal}}$.
G.1 Synthetic feature and task generation
All stress tests use a controlled synthetic ICL surrogate where the only changing factor is the feature distribution. We generate a sequence of feature vectors $\{z_{t}\}_{t=1}^{K}\subset\mathbb{R}^{d}$ (the “demonstration representations”) and define a simple supervised task by a random linear teacher. Specifically, we draw a teacher vector $w^{\star}\sim\mathcal{N}(0,I_{d})$ and assign labels $y_{t}=\mathrm{sign}(\langle w^{\star},z_{t}\rangle+\eta_{t})$ with i.i.d. noise $\eta_{t}\sim\mathcal{N}(0,\sigma^{2})$. Given $K$ demonstrations, the in-context predictor is instantiated by ridge regression on the synthetic features:
$\widehat{w}_{K}\in\arg\min_{w}\sum_{t=1}^{K}(y_{t}-\langle w,z_{t}\rangle)^{2}+\lambda\|w\|_{2}^{2},\qquad\widehat{y}(x)=\mathrm{sign}(\langle\widehat{w}_{K},z(x)\rangle).$
We evaluate $\mathrm{ACC}(K)$ on a held-out synthetic test set generated from the same setting and estimate $K_{\text{emp}}$ using the same knee-point estimator as in Appendix C. Uncertainty in $K_{\text{emp}}$ is quantified by percentile bootstrap over test instances (with $B=200$ resamples, matching Appendix C).
Nominal baseline.
In the nominal setting (A1), we sample $z_{t}\sim\mathcal{N}(0,\Sigma)$ i.i.d. with a fixed PSD matrix $\Sigma$.
Bounded drift (A2).
To model bounded representation drift, we sample i.i.d. base features $\tilde{z}_{t}\sim\mathcal{N}(0,\Sigma)$ and add a bounded perturbation
$z_{t}=\tilde{z}_{t}+\Delta_{t},\qquad\|\Delta_{t}\|_{2}\leq\varepsilon.$
Unless stated otherwise, we set $\Delta_{t}=\varepsilon\,u$ with a fixed unit vector $u$ (systematic drift), which makes the violation controlled and reproducible; random-direction drift can be generated by sampling $u_{t}$ uniformly on the unit sphere.
Heavy tails (A3).
To model heavy-tailed features, we sample $z_{t}$ from a multivariate Student-$t$ distribution with $\nu$ degrees of freedom and rescale it to match a target second-moment:
$z_{t}=\sqrt{\frac{\nu-2}{\nu}}\,\Sigma^{1/2}\,g_{t},\qquad g_{t}\sim t_{\nu}(0,I_{d}),\quad(\nu>2).$
Lower $\nu$ yields heavier tails while preserving the same second-moment scale.
Temporal dependence (A4).
To introduce dependence while keeping the same marginal second-moment, we use a stationary AR(1) process:
$z_{t}=\rho\,z_{t-1}+\sqrt{1-\rho^{2}}\,\varepsilon_{t},\qquad\varepsilon_{t}\sim\mathcal{N}(0,\Sigma),$
so that $\mathbb{E}[z_{t}z_{t}^{\top}]=\Sigma$ for all $t$ and the correlation strength increases with $\rho\in[0,1)$.
G.2 Bounded Drift
@@
G.3 Heavy-tailed Features
@@
Heavier tails inflate the required prompt length by a multiplicative constant, but truncation or normalization substantially reduces the inflation. This behavior aligns with the extended bound in Appendix F.
G.4 Temporal Dependence
@@
Empirical knee-points scale approximately as $(1+c_{\rho})\cdot K_{\mathrm{nominal}}$, consistent with our dependence extension in Appendix E.
G.5 Summary
@@
Appendix H Interpretability of the Proxy
Proposition H.1 (Predictive variance control via $\lambda_{\min}$).
@@ Appendix H Interpretability of the Proxy
See H.1 for details; it follows by the matrix inequality $A^{-1}(K\widehat{\Sigma}_{K})A^{-1}\preceq A^{-1}$ with $A=K\widehat{\Sigma}_{K}+\lambda I$ and the Rayleigh quotient bound $\mathbf{v}^{\top}A^{-1}\mathbf{v}\leq\|\mathbf{v}\|_{2}^{2}/\lambda_{\min}(A)$. ∎
@@
Proposition H.1 (Appendix H) shows that a lower bound on $\lambda_{\min}(\hat{\Sigma}_{K})$ suffices to control predictive variance of in-context regression, providing a mechanistic rationale for why it correlates with stability.
@@
H.1 Proof sketch of Proposition
@@ Conclusion.
which establishes Proposition H.1. $\square$
@@
Appendix I Ablation Studies
@@
To bridge the gap between theory and practice, we performed ten ablations that correspond to the key factors appearing in our formula and experimental pipeline. We first present a summary table that aggregates all results, followed by ten dedicated “subsections, each restating its own result with brief analysis and takeaway. All numbers are averages over multiple runs unless otherwise noted, and error ratios are defined as theoretical $K^{*}$ divided by the empirical knee (lower is better).
I.1 $\delta$-ratio Sensitivity
@@
I.2 Effective Rank Definition
@@ I.3 Sub-Gaussian Parameter $\sigma$
I.4 Second-moment Norm $\|\Sigma\|$
@@ I.5 Label Tokenization
I.6 Prompt Truncation Policy
@@
I.7 Theory Constant Scaling
@@
I.8 Synthetic Data Experiments
@@ I.9 Retrieval Strategy
Interpretation. Retrieval reshapes the empirical second-moment around the test point, preserving a larger effective $\lambda_{\min}$ and tightening $K^{*}$.
@@ I.10 Model Scale
Appendix J Theoretical Constants: Sources, Mappings, and Practical Calibration
J.1 List of Constants and Quick Reference
@@
Symbol Type Occurrence Recommended Range
2, 2/32,\,2/3 Structural Matrix Bernstein tail bound Fixed, non-tunable
$c_{1}$ Absolute Fourth-moment bound $[1.5,6]$, recommend 3344
$c_{2}$ Absolute Second-moment deviation bound $[2,4]$, recommend 33
$c^{\prime},\,c_{3}$ Engineering Sample-size formulae $\geq 8$, typically 10101515
J.2 Theoretical Origins and Geometric Interpretation
@@
For the sum of independent centered random matrices $S=\sum Y_{k}$, then
@@
second-moment operator deviation constant $c_{2}$.
Koltchinskii & Lounici proved
@@
J.3 Coupling with Problem Parameters
@@
J.4 Three-Step Calibration Procedure
@@
Choosing the target floor $\delta$ and handling infeasible targets.
The proxy objective in (2) is parameterized by a target spectral floor $\delta$, which controls how well-conditioned the demonstration second moment must be to certify the surrogate stability event. In Algorithm 1, the check $\widehat{\Delta}=\underline{\lambda}-\delta>0$ is necessary; if the algorithm returns infeasible target, it means that the requested $\delta$ is larger than the estimated population floor at the desired confidence level, under the current feature map and ridge level. In practice we recommend selecting $\delta$ by validation: sweep $\delta$ on a small held-out set, choose the smallest value that achieves the desired stability/accuracy trade-off, and then lock it for test-time estimation. When available, Appendix L provides a direct estimator for distributional stability in the sense of Definition A.1 that can serve as the validation signal. If infeasibility persists, reduce $\delta$, increase the ridge via $\epsilon_{\rho}$, or use the quantile-floor variant described in Section 3.5.
Appendix K Multi-step Reasoning and RAG Tasks
This appendix specifies the experimental protocol for multi-step reasoning and retrieval-augmented generation (RAG). Unless stated otherwise, knee estimation, bootstrap CIs, and calibration follow Appendix C.
K.1 Multi-step reasoning (GSM8K / SVAMP)
Prompt template (final answer only).
We constrain outputs to a canonical final-answer format to reduce ambiguity while preserving multi-step computation. Each prompt contains $K$ demonstrations followed by a query:
Task: Grade-school math word problems.
Solve the problem and output only the final numeric answer.
Q: {q_1}
A: {a_1}
Q: {q_K}
A: {a_K}
Q: {q_q}
A:
Example (one demo + one query).
Task: Grade-school math word problems.
Solve the problem and output only the final numeric answer.
Q: Lily has 12 candies. She gives 5 to her friend. How many candies does she have left?
A: 7
Q: A store sold 18 apples in the morning and 27 in the afternoon. How many apples were sold in total?
A:
Evaluation.
We normalize the produced answer string (strip spaces/punctuation) and compute exact match against the gold final answer.
$K$ grid.
Due to longer outputs, we optionally use a cost-aware grid $\{0,1,2,4,8,16,32,64\}$ (or a prefix thereof).
K.2 RAG (HotpotQA / NaturalQuestions)
Retrieval.
We retrieve top-$m$ evidence passages per query (default $m=5$) from a fixed corpus snapshot using a deterministic retriever (e.g., BM25). The retrieved set for each query is held fixed across all $K$.
Prompt template (evidence before demos).
We place evidence before demonstrations to reduce interleaving artifacts:
Task: Answer the question using the evidence. Output only the short final answer.
Evidence:
[1] {p_1}
[m] {p_m}
Q: {q_1}
A: {a_1}
Q: {q_K}
A: {a_K}
Q: {q_q}
A:
Example (evidence + one demo + one query).
Task: Answer the question using the evidence. Output only the short final answer.
Evidence:
[1] Marie Curie was a physicist and chemist who conducted pioneering research on radioactivity.
[2] She won the Nobel Prize in Physics in 1903 and the Nobel Prize in Chemistry in 1911.
Q: How many Nobel Prizes did Marie Curie win?
A: 2
Q: In what year did Marie Curie win the Nobel Prize in Chemistry?
A:
Evaluation.
We score exact match and token-level F1 using the dataset’s standard evaluation protocol with standard normalization.
Spectral proxy under RAG.
By default, representations $\phi(x)$ are computed from the question text only to avoid retrieved evidence dominating second-moment estimates. As an ablation, we compute $\phi(x)$ from question+evidence and report its impact.
K.3 Extended results (reference values)
Table 10: Extended error ratios on additional reasoning/RAG datasets and RAG representation ablations. Lower is better.
Dataset / Variant Raw (mean) Calibrated (mean)
[…]
Analysis.
Multi-step reasoning and RAG exhibit systematically larger raw error ratios than classification, consistent with higher demo sensitivity and additional prompt interactions. Calibration reduces residual conservativeness in both families, typically bringing ratios close to $1\times$. For RAG, using question-only $\phi$ tends to provide a cleaner second-moment signal; incorporating evidence can inflate the proxy and degrade tightness when evidence is noisy or heterogeneous.
Appendix L Direct Measurement of Distributional Stability (Definition 3.1)
This appendix directly measures the distributional stability criterion introduced in Definition 3.1, rather than relying on accuracy-based proxies. We explicitly quantify the variability of the model’s output distribution under independent resampling of demonstrations, and study how the resulting stability thresholds align with (i) empirical accuracy knee-points and (ii) our spectral proxy.
L.1 Output distribution and divergence metric
For a fixed query $x_{q}$ and a sampled demonstration set $S_{K}$ of size $K$, the prompted model induces a distribution over candidate answers,
$p_{S_{K}}(y\mid x_{q}),\qquad y\in\mathcal{Y}.$
For tasks with a closed answer set (e.g., multiple-choice), $\mathcal{Y}$ is the set of options. For open-ended tasks, $\mathcal{Y}$ is defined as the union of the top-$M$ decoded candidates across all seeds and $K$ values, with $M=20$; probabilities are renormalized on this support.
To compare two output distributions, we use Jensen–Shannon divergence (JSD),
$\mathrm{JSD}(p,q)=\frac{1}{2}\mathrm{KL}(p\|m)+\frac{1}{2}\mathrm{KL}(q\|m),\qquad m=\tfrac{1}{2}(p+q),$
measured in nats. JSD is symmetric, bounded, and empirically stable for finite-support distributions.
L.2 Estimating stability probability
Definition 3.1 requires that, with high probability over independently sampled demonstration sets, the divergence between the resulting output distributions remains below a tolerance $\tau$.
For each query $x_{q}$, demonstration size $K$, and tolerance $\tau$, we estimate this probability by repeated resampling. Specifically, we draw $R=50$ independent pairs of demonstration sets $\{(S_{K}^{(r)},S_{K}^{\prime(r)})\}_{r=1}^{R}$ and compute
$d_{q}^{(r)}(K)=\mathrm{JSD}\!\left(p_{S_{K}^{(r)}}(\cdot\mid x_{q}),p_{S_{K}^{\prime(r)}}(\cdot\mid x_{q})\right).$
The per-query stability estimate is
$\widehat{\pi}_{q}(K;\tau)=\frac{1}{R}\sum_{r=1}^{R}\mathbf{1}\!\left[d_{q}^{(r)}(K)\leq\tau\right].$
Aggregating over a dataset of $Q=200$ queries, we report the empirical stability probability
$\widehat{\Pi}(K;\tau)=\frac{1}{Q}\sum_{q=1}^{Q}\widehat{\pi}_{q}(K;\tau).$
For a fixed failure tolerance $\xi$, we define the direct stability threshold
$K^{\star}_{\tau,\xi}=\min\left\{K:\widehat{\Pi}(K;\tau)\geq 1-\xi\right\}.$
Unless stated otherwise, we use $\xi=0.10$ and $\tau\in\{0.05,0.10\}$.
L.3 Alignment with knee-points and spectral proxy
Table 11 compares the directly measured stability threshold $K^{\star}_{\tau,0.10}$ with the accuracy knee-point $K_{\mathrm{knee}}$ and the theory-driven spectral proxy prediction $K_{\mathrm{proxy}}$. Across all evaluated tasks, the three quantities are closely aligned, with discrepancies typically within a small constant.
Task $\tau$ $K^{\star}_{\tau,0.10}$ (direct) $K_{\mathrm{knee}}$ $K_{\mathrm{proxy}}$ $|K^{\star}-K_{\mathrm{knee}}|$ $|K^{\star}-K_{\mathrm{proxy}}|$
MC Reasoning 0.05 16 16 16 0 0
MC Reasoning 0.10 8 8 8 0 0
Fact Retrieval 0.05 24 24 24 0 0
Fact Retrieval 0.10 16 16 16 0 0
Open Generation 0.05 32 28 32 4 0
Open Generation 0.10 16 16 16 0 0
Table 11: Direct measurement of distributional stability (Definition 3.1) using Jensen–Shannon divergence. $K^{\star}_{\tau,0.10}$ denotes the smallest $K$ such that the empirical stability probability exceeds $0.9$. Results are compared with the accuracy knee-point and the spectral proxy prediction.
L.4 Stability probability and divergence quantiles
To further characterize the stability transition, Table 12 reports per-query quantiles of JSD aggregated over resampled demonstration pairs, together with the empirical stability probability $\widehat{\Pi}(K;\tau=0.05)$. As $K$ increases, both the median and upper-tail divergences decrease sharply, and the $90\%$ stability criterion is reached near the predicted threshold.
Task $K$ Q50 Q75 Q90 Q95 $\widehat{\Pi}(K;\tau=0.05)$
MC Reasoning 8 0.044 0.062 0.087 0.104 0.63
MC Reasoning 16 0.019 0.029 0.042 0.053 0.91
MC Reasoning 24 0.012 0.019 0.028 0.036 0.96
Fact Retrieval 16 0.041 0.058 0.084 0.102 0.67
Fact Retrieval 24 0.020 0.031 0.046 0.059 0.90
Fact Retrieval 32 0.014 0.022 0.033 0.042 0.95
Table 12: Quantiles of Jensen–Shannon divergence across queries and resampled demonstration pairs. The collapse of the upper tail coincides with the stability threshold predicted by both the knee-point and the spectral proxy.
L.5 Discussion: consistency and boundary cases
These results confirm that the stability transition predicted by Definition 3.1 is directly observable at the distributional level and occurs at nearly the same prompt length as the accuracy knee-point and the spectral proxy threshold. When minor discrepancies arise, they are attributable to principled effects rather than noise.
First, in some open-generation settings, top-1 accuracy saturates slightly earlier than full distributional stabilization, leading to $K^{\star}_{\tau,\xi}>K_{\mathrm{knee}}$ for small $\tau$. Second, distributional stability does not imply correctness: a model may consistently assign high probability to an incorrect answer across resampled demonstrations, yielding low divergence but poor accuracy on a subset of ambiguous queries. Finally, the spectral proxy may overestimate the required $K$ when the external encoder geometry overstates the diversity of demonstrations that the model effectively attends to.
Overall, directly measuring distributional stability closes the gap between theory and experiment by validating the stability criterion itself, rather than relying solely on accuracy-based surrogates.